PartnerObserving Log for 2007 June 20-21

Observer: D. Thompson
LBC Support: A. DiPaola, R. Speziali, R. Bertram
Telescope Operator: D. Gonzalez

23:00 AdP & RS found the port server 12V power supply at 11.2V. They trimmed it up to 12.2V and housekeeping came up without errors. They believe it is fixed.

23:03 Ran 5bias to see if everything is working. Data not appearing in the /Repository.

Workaround (from lbtmu03 xterm): cd /Repository
sftp root@lbcarchive2 (passwd in black LBC notebook)
cd /data/store/indir
get lbcb.20070620... (replace with filename)

02:40 Sunset: TCS and LBC are ready to go. Crew still working on inserting a pivot pin in the M3 swingarm. We cannot move the telescope until they are done.

03:18 Telescope back together (yeah!), cleanup and checkout.

03:33 25bias at horizon while David starts things up.

03:43 Opening dome

03:55 Ready. Tout=14.5, Tmirror=10.2

03:56 Sent SA107_Vgr, LBC is not receiving Az, El, Airmass from TCS. LBC interface flashing repeatedly, but it recovered after a few minutes. Preset was to dec=-00:13:26, but it looks like it went to +00... Also unable to send -00 dec from preset GUI, -01 worked. Michele called. Chris called. [Note added later by JMH: Michele found and corrected a bug in the PCS GUI related to the display of -00 Dec coordinates. The telescope was going to the right place, but the display was wrong.]

04:35 AdP fixed Az, El, Airmass problem - keywords changed on TCS side and he was not notified. LBC now seems to be running fine.

05:42 Michele reports most of the multiple problems fixed. LSS disabled.

05:44 Vfastextra @ 17h +20d, Airmass 1.037 054631 Shutter did not open? Image looks like a bias! AdP cycled power just to the shutter controller from lbccontrol eng. 055135 also bad 055440 also bad

AdP suggests rotating to PA=180 and try again... 060011 nope... 060358 and again nope...

06:07 Closing dome so AdP & RS can go have a look.

### Note: Moonset 23:53 MST 17:23 LST ###

064803 RS reports shutter telemetry and command wires appear to be broken. The camera will need to be dismounted to be worked on in the lab downstairs. There is also a rotator imit switch that is "destroyed". It is possible that these are related. RS & AdP prefer to do the repairs tomorrow so they can also replace some dying fans and the limit switch. So...we are toast for the night. Shutting down.

[Note added later by JMH: The shutter cable was repaired the next day by Roberto S. The limit switch still needs to be replaced. See Issue #704.]

06:56 Done for the night. 100% loss to technical problems on both the telescope (both hardware and software gave problems) and LBC (shutter failure).

-- JohnHill - 22 Jun 2007 (from D. Thompson's notes.
Topic revision: r4 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback